# Hidden memory and stochastic fluctuations in science

### v1

@@ VI.1 Philosophical and theoretical implications
Furthermore, these findings hold significant implications for real-world applications of fractional Brownian motion. It is well known that its mathematical structure fundamentally differs depending on whether the Hurst parameter is greater or less than $\tfrac{1}{2}$ [55, 76]. Historically, the regime $H\in(\tfrac{1}{2},1)$ has been more extensively studied, particularly in finance, due to the long-range dependence observed in certain market data. In contrast, the case $H\in(0,\tfrac{1}{2})$, where citation dynamics fall, has received comparatively less attention in real-world applications. However, models exhibiting anti-persistence in this regime have recently gained significant attention in finance, particularly in the study of rough volatility [65].
The theoretical framework developed in this study applies broadly to both anti-persistent and persistent systems. Given the universality of the adopted Ansatze (i)(iv), it extends beyond conventional citation processes and may serve as a foundational theory for analogous attention dynamics [77, *Lorenz-Spreen19]. In the empirical analysis in Section V, we examined citation as a well-defined example of attention, focusing on annual citation counts. If the theory applies more broadly, other forms of attention should also adhere to the $t^{H}$ law when the SD of the logarithm of attention at each point are analysed. Moreover, cumulative attention is generally well approximated by a log-normal distribution but may display power-law behaviour depending on volatility, memory effects, and the shape of the average attention history curve [53]. These insights indicate diverse applications across multiple fields, highlighting the need for further research to refine and validate the theory’s broader relevance.
@@ VI.2 Memory effects and non-Markovianity
[…]
@@ VI.3 Unanswered questions and conjecture
One possible resolution to these questions is suggested by classical mathematical arguments based on Donsker’s invariance principle [82, *Taqqu75, *Enriquez04]. Drawing on this theory, we propose the following conjecture: Beneath citation processes, which are inherently discrete, there exists an underlying integer-valued stochastic process. By repeating this process many times and applying appropriate normalisation, the functional central limit theorem ensures the emergence of a discrete centred Gaussian process. Taking the scaling limit of this discrete process then naturally leads to fractional Brownian motion as a characteristic feature of the system.
@@ VII Summary and Conclusion
Specifically, we adopted Ansatze (i)(iv) (Section II) and formulated the fractional SDE (8) that governs latent attention, whose discrete realisation corresponds to citation counts. By linking its solution, Eq. (9), to citation dynamics via Eq. (7), we demonstrated—both theoretically (Section III) and through numerical simulations (Section IV)—that the resulting model satisfies four key properties:
@@
Furthermore, applying our validated model to empirical data on arXiv e-prints has enhanced the microscopic understanding of citation phenomena observed in real-world settings (Section V). Additionally, we have successfully estimated the fractal dimension associated with the system’s attention.
Thus, the in-degree distribution of citation networks, a long-standing subject of research, is naturally explained within our theoretical framework based on a few key system characteristics: the Hurst parameter ($H$), the volatility parameter ($\beta$), and the time dependence of the deterministic component ($\alpha(t)$ or $A(t)$). A central aspect of this model is the use of fractional Brownian motion to characterise the system’s stochastic component, with the memory effect governed by the Hurst parameter $H$.
Specifically, we demonstrated that when $H\in(0,\tfrac{1}{2})$, indicating an anti-persistent nature, and this is combined with the empirical observation that the average citation history curve decays over time, the model effectively accounts for why citation distributions are often well approximated by log-normal distributions. Conversely, when fluctuations occasionally lead to sustained high levels of attention, the citation distribution is better approximated by a power law. This distinction between the two regimes offers a coherent explanation for the empirical variability observed in citation dynamics.
As discussed in Section VI, this perspective also gives rise to new questions from physical, mathematical, and statistical viewpoints. Addressing these open problems is expected to further advance our understanding of network dynamics and foster future developments in nonlinear dynamics, non-equilibrium statistical mechanics, as well as chaos and fractal theory.
@@ Acknowledgements.
The views and conclusions expressed herein are solely those of the author and should not be construed as necessarily reflecting the official policies or endorsements—whether explicit or implicit—of any organisations with which the author is presently or has previously been affiliated.
@@
for ‘Hidden memory and stochastic fluctuations in science’ by K. Okamura (2025).
This Supplementary Materials file includes:
  • Supplementary Text for Sections IIVI
  • Figures S2S6
@@
The $t^{H}$ law (1) discussed in the main text could, in principle, be reproduced within the stochastic model introduced in [23], which extends the Black–Scholes model by incorporating a time-dependent volatility function. This can be achieved by appropriately tuning the volatility function to align with the $t^{H}$ law. Under this interpretation, the time dependence of volatility may be attributed to factors such as improved discoverability due to technological advancements or, conversely, reduced visibility caused by the rapid growth in publication volume. In this sense, the model provides a potential explanation for the observed phenomenon.
@@
Supplementary Text for Section III
@@
There are several ways to describe a fractional Brownian motion, $\{B_{H}(t)\}_{t\in\mathbb{R}}$, with Hurst parameter $H\in(0,1)$, in terms of the standard Brownian motion, $\{W(t)\}_{t\in\mathbb{R}}$; see [76, Section 1.2] and the references therein. Among these approaches, we introduce one that employs an operator $M_{H}$ acting on a rapidly decreasing smooth function $f$ defined on $\mathbb{R}$, explicitly given by [67]
[…]
where $\Gamma(\cdot)$ represents the gamma function. Choose $f(s)$ as the indicator function $\chi_{[a,b]}(s)$, which equals $1$ if $a\leq s\leq b$, $-1$ if $b\leq s\leq a$ (except when $s=a=b$), and 00 otherwise. We then integrate Eq. (S1) over the entire real line with respect to $\mathop{}\!dW(s)$. Setting $a=0$ and $b=t$, we obtain
[…]
where $\Theta(\cdot)$ denotes the Heaviside step function, defined as $\Theta(x)=1$ for $x>0$ and $\Theta(x)=0$ for $x\leq 0$. Equation (S2) matches, up to a normalisation factor, the original expression for $B_{H}(t)$ introduced by Mandelbrot and Van Ness [55, Eq. (2.1)].
@@
where the operator $M_{H}(\cdot)$ and the indicator function $\chi_{[a,b]}(\cdot)$ are as defined earlier. In the main text, we considered the special case in which $\beta(s)=\beta$ is constant, and under that condition, the solution (S4) reduces to Eqs. (9, 10).
Supplementary Text for Section IV
[…]
In the main text, when simulating $C(T)$ as expressed in Eq. (7) based on our theoretical model, we introduced the approximation $\hat{C}(T)$ as given in Eq. (17). This approximation always yields a natural number for $T>0$, even in cases where $X_{i}\mathop{}\!\Delta t$ remains negligibly small throughout $i=1,\,\dots,\,n$, such that its summation over the entire period does not even reach 1. Nonetheless, Eq. (17) ensures that $\hat{C}=1$ even in such situations. In this respect, the effect is analogous to the ‘$\text{citation}+1$’ prescription often used in the literature [35, 32, 24, 23], which facilitates the computation of the logarithm of citation counts even in cases with zero citations. We assume that the bias [70] introduced by employing the ceiling function here is negligible and does not affect our main conclusions.
@@
Figure S2 shows simulated paths of $X(t)$ under $H=0.85$ and $\alpha(t)=1$. In contrast to the anti-persistent case (Fig. 3), the trajectories here are notably smoother, with fewer pronounced fluctuations. In Fig. S2(c), $\{X_{i}\}_{i=1}^{n}$ follows a nearly steady trend and eventually declines to zero at time $T$. However, this outcome is not inevitable. Due to the persistent nature of the process, $\{X_{i}\}_{i=1}^{n}$ often continues along a certain trajectory, yet it could also move in the opposite direction and increase indefinitely. Such paths contribute to large values of $\hat{C}(T)$.
Figure S2 shows the results of Monte Carlo simulations with $N_{\mathrm{s}}=50{,}000$, conducted under the same conditions. Additionally, Fig. S3(a) presents the kernel density estimates of the distribution for $\beta=0.2$, $0.5$, $0.8$, $1.0$ and $1.5$. Compared with the anti-persistent case (Fig. 3), the distribution of $C(T)$ now has a substantially fatter tail, a tendency that grows more pronounced with increasing $\beta$.
Moreover, the Q–Q plot in Fig. S3(b) deviates noticeably from a straight line, indicating that the distribution of $\hat{C}(T)$ no longer closely follows a log-normal form. This does not necessarily mean that a power law is the best fit. However, as shown in Fig. S3(c) and Table S1, where the upper 1% of $\hat{C}(T)$ is fitted with a shifted power law, the persistent scenario exhibits a stronger tendency to be well approximated by a shifted power law than the anti-persistent case. Still, at extremely high citation counts, the data also deviate from the shifted power law.
Supplementary Text for Section V
@@
The data used in Fig. 1 and the empirical data analysis in Section V of the main text are the same as those employed in [23], with the raw data compiled in [54]. For further details on the data, please refer to these references.
@@
In the main text, the prospective approach is used (Section II.2). However, using the retrospective approach yields qualitatively similar results, again confirming the $t^{H}$ law trend. Indeed, when the $t^{H}$ law was first observed, the retrospective approach was applied [23, Appendix C, Suppl. Fig. S9].
@@
In the main text, the estimation of the Hurst parameter $H$ and volatility $\beta$ was carried out on the entire aggregate set of e-prints across six arXiv disciplines—Astrophysics (‘astro-ph’), Computer Science (‘comp-sci’), Condensed Matter Physics (‘cond-mat’), High Energy Physics (‘hep’), Mathematics (‘math’) and other physics (‘oth-phys’)—the same classification as in [23, 54]. Figure S6 shows the results of the same analysis, now broken down by discipline. From Fig. S6(a), it is clear that the memory effect is anti-persistent in all fields, but a consistent historical trend indicates that High Energy Physics has the strongest anti-persistent tendency, whereas Condensed Matter Physics has the weakest. Additionally, Fig. S6(b) reveals that volatility is especially high in High Energy Physics and Computer Science.
These observations likely reflect each discipline’s distinct characteristics, such as scientific discovery patterns, citation behaviours in research communities, and publishing norms—factors also discussed in the main text. Combined with earlier analyses of discipline-specific spatiotemporal aspects of citation evolution and with measures such as the ‘internal obsolescence rate’ and ‘retention rate’ reported in [23], these results provide useful insights not only from scientometric or bibliometric standpoints, but also within the broader ‘science of science’ framework [75].
@@
Typically, the Hurst parameter for fractional Brownian motion is estimated via methods such as the Rescaled Range Analysis (R/S) method [85, 72], or improved variants thereof. This approach involves calculating the range of partial sums of deviations for time series segments from their means, then normalising by the corresponding SD. However, it is not suitable in te current study. The reason is that the presence of an unknown, time-dependent deterministic nonlinear term $\alpha(t)$ disrupts self-similar scaling properties, thereby biasing R/S-based estimates.
Moreover, the arXiv dataset used in this study [54] records citation counts on a yearly basis, yielding at most a few dozen data points per e-print. Methods like R/S analysis or other estimation techniques commonly employed in financial engineering [76, 56] usually require much longer time series. Despite recent advances in bibliometric data, collecting extended citation histories at finer temporal resolutions (e.g. monthly or daily) remains practically challenging. Where monthly-level citation records exist, they often involve data imputation: for instance, if the exact publication month is missing, the record may default to January. Similarly, daily-level data can default to the first of the month if only the month is known. Hence, careful preprocessing is needed to handle missing values or exclude certain entries, making the construction of precise time series difficult at monthly or daily resolutions.
Given these constraints, this study employed a heuristic estimation of the Hurst parameter based on the newly discovered $t^{H}$ law. Although the resulting estimates are approximate, their qualitative features—such as the evidence of anti-persistent behaviour—are statistically robust within the limits of the applied tests.
@@
Now, consider a hypothetical world in which all publishing constraints—whether resource-based, technological, or conventional—are entirely removed. In this setting, researchers would have optimal access to literature search, knowledge acquisition, and reference management. Citing older publications would be actively encouraged, with no practical limit on the number of references included. Under such conditions, the typical citation history curve, when aggregated across numerous publications, might no longer follow a jump-and-decay pattern but instead develop a heavy-tailed form over time or even evolve into a monotonically increasing function.
In this scenario, the distribution of the time-integrated attention measure would likely be better represented by a power law or another even heavier-tailed distribution, rather than a log-normal distribution. Alternatively, it is possible that such an attention-dynamics paradigm already exists in certain domains, even if it has not yet been widely recognised.
@@
Figure S4: Relationship between the standard deviation (SD) of the logarithm of citation counts acquired each year up to February 2020 and the number of years ($t$) elapsed since posting, shown for e-prints published on arXiv between 1995 and 2006. The data used are the same as in [54].
Figure S5: Estimated values of (a) $\hat{H}$ and (b) $\hat{\beta}$, along with their standard errors for e-prints published on arXiv from 1995 to 2006, presented for each publication year (numerical data are provided in Table S2). The data used are the same as in [54].
Figure S6: Estimated values of (a) $\hat{H}$ and (b) $\hat{\beta}$ for e-prints published on arXiv from 1999 to 2010, shown for six different arXiv disciplines—Astrophysics (‘astro-ph’), Computer Science (‘comp-sci’), Condensed Matter Physics (‘cond-mat’), High Energy Physics (‘hep’), Mathematics (‘math’) and other physics (‘oth-phys’). The data used are the same as in [54].

### v2

@@ VI.1 Philosophical and theoretical implications
Furthermore, these findings hold significant implications for real-world applications of fractional Brownian motion. It is well known that its mathematical structure fundamentally differs depending on whether the Hurst parameter is greater or less than $\tfrac{1}{2}$ [56, 81]. Historically, the regime $H\in(\tfrac{1}{2},1)$ has been more extensively studied, particularly in finance, due to the long-range dependence observed in certain market data. In contrast, the case $H\in(0,\tfrac{1}{2})$, where citation dynamics fall, has received comparatively less attention in real-world applications. However, models exhibiting antipersistence in this regime have recently gained significant attention in finance, particularly in the study of rough volatility [66].
The theoretical framework developed in this study applies broadly to both antipersistent and persistent systems. Given the universality of the adopted Ansätze (i)(iv), it extends beyond conventional citation processes and may serve as a foundational theory for analogous attention dynamics [82, *Lorenz-Spreen19]. In the empirical analysis in Sec. V, we examined citation as a well-defined example of attention, focusing on annual citation counts. If the theory applies more broadly, then other forms of attention should also adhere to the $t^{H}$ law when the SD of the logarithm of attention at each point are analysed. Moreover, cumulative attention is generally well approximated by a log-normal distribution but may display power-law behaviour depending on volatility, memory effects, and the shape of the average attention history curve [53]. These insights indicate diverse applications across multiple fields, highlighting the need for further research to refine and validate the theory’s broader relevance.
@@ VI.2 Memory effects and non-Markovianity
Further progress may also be achieved by refining how stochastic fluctuations are introduced. Potential directions include employing mixed fractional Brownian motions [84, *Thale09] or reset geometric fractional Brownian motion [86]. For instance, given the interdisciplinary variations in the estimated $\hat{H}$ and $\hat{\beta}$ values observed in Sec. V (Fig. S6), it is plausible that citation dynamics across the broader academic landscape, including the arXiv disciplines used here as a testbed, might be better described by a mixed fractional Brownian motion. Such a process can be expressed as either a discrete sum, $Z(t)=\sum_{\ell\in\mathcal{S}}\beta_{\ell}B_{H_{\ell}}(t)$, or a continuous analogue, $Z(t)=\int_{\mathcal{S}}\beta(\rho)B_{H(\rho)}(t)\mathop{}\!d\rho$. Here $B_{H_{\ell}}(t)$ are independent fractional Brownian motions with $H_{\ell}\in(0,1)$ and $\beta_{\ell}>0$, where $\mathcal{S}$ denotes a proper summation domain, with a similar formulation applying to the continuous version.
@@ VI.3 Unanswered questions and conjecture
One possible resolution to these questions is suggested by classical mathematical arguments based on Donsker’s invariance principle [87, *Taqqu75, *Enriquez04]. Drawing on this theory, we propose the following conjecture: beneath citation processes, which are inherently discrete, there exists an underlying integer-valued stochastic process. By repeating this process many times and applying appropriate normalisation, the functional central limit theorem ensures the emergence of a discrete centred Gaussian process. Taking the scaling limit of this discrete process then naturally leads to fractional Brownian motion as a characteristic feature of the system.
@@ VII Summary and Conclusion
Specifically, we adopted Ansätze (i)(iv) (Sec. II) and formulated the fractional SDE (8) that governs latent attention, whose discrete realisation corresponds to citation counts. By linking its solution, Eq. (9), to citation dynamics via Eq. (7), we demonstrated—both theoretically (Sec. III) and through numerical simulations (Sec. IV)—that the resulting model satisfies four key properties:
@@
Furthermore, applying our validated model to empirical data on arXiv e-prints has enhanced the microscopic understanding of citation phenomena observed in real-world settings (Sec. V). Additionally, we have successfully estimated the fractal dimension associated with the system’s attention.
Thus, the in-degree distribution of citation networks, a long-standing subject of research, is naturally explained within our theoretical framework based on a few key system characteristics: the Hurst parameter ($H$), the volatility parameter ($\beta$), and the time dependence of the deterministic component $\alpha(t)$ or $A(t)$]. A central aspect of this model is the use of fractional Brownian motion to characterise the system’s stochastic component, with the memory effect governed by the Hurst parameter $H$.
Specifically, we demonstrated that when $H\in(0,\tfrac{1}{2})$, indicating an antipersistent nature, and this is combined with the empirical observation that the average citation history curve decays over time, the model effectively accounts for why citation distributions are often well approximated by log-normal distributions. Conversely, when fluctuations occasionally lead to sustained high levels of attention, the citation distribution is better approximated by a power law. This distinction between the two regimes offers a coherent explanation for the empirical variability observed in citation dynamics.
As discussed in Sec. VI, this perspective also gives rise to new questions from physical, mathematical, and statistical viewpoints. Addressing these open problems is expected to further advance our understanding of network dynamics and foster future developments in nonlinear dynamics, nonequilibrium statistical mechanics, as well as chaos and fractal theory.
@@ Acknowledgements.
The author would like to thank the anonymous referees of Phys. Rev. E for their valuable comments on the manuscript. The views and conclusions expressed herein are solely those of the author and should not be construed as necessarily reflecting the official policies or endorsements—whether explicit or implicit—of any organisations with which the author is presently or has previously been affiliated.
Data Availability
The data that support the findings of this article are openly available [55, 90].
@@
for ‘Fractional stochastic model of citation dynamics with memory and volatility’
This Supplemental Material accompanies the article published in Phys. Rev. E 112, 044304 (2025), DOI: 10.1103/l2xd-43n9.
It contains:
  • Supplementary text for Secs. IIVI
  • Figs. S2S6
@@
The $t^{H}$ law (1) discussed in the main text could, in principle, be reproduced within the stochastic model introduced in Ref. [23], which extends the Black–Scholes model by incorporating a time-dependent volatility function. This can be achieved by appropriately tuning the volatility function to align with the $t^{H}$ law. Under this interpretation, the time dependence of volatility may be attributed to factors such as improved discoverability due to technological advancements or, conversely, reduced visibility caused by the rapid growth in publication volume. In this sense, the model provides a potential explanation for the observed phenomenon.
@@
Note also that the empirically estimated upper bound of the Hurst exponent for the arXiv data, as reported in Sec. V of the main text, is relatively small, around $H\approx 0.13$. When $H$ is sufficiently small, the power-law growth $t^{H}$ can be approximated by $1+H\ln t$. This approximation helps explain the near-logarithmic behaviour observed in the temporal evolution of citation variance, as previously illustrated in Ref. [23, Appendix C, Suppl. Fig. S9].
Supplementary Text for Sec. III
@@
There are several ways to describe a fractional Brownian motion, $\{B_{H}(t)\}_{t\in\mathbb{R}}$, with Hurst parameter $H\in(0,1)$, in terms of the standard Brownian motion, $\{W(t)\}_{t\in\mathbb{R}}$; see Ref. [81, Sec. 1.2] and the references therein. Among these approaches, we introduce one that employs an operator $M_{H}$ acting on a rapidly decreasing smooth function $f$ defined on $\mathbb{R}$, explicitly given by [68]
$M_{H}f(s)=\begin{dcases*}\,C_{H}\int_{\mathbb{R}}\left|u\right|^{H-\frac{3}{2}}\big[f(s-u)-f(s)\big]\mathop{}\!du&for $H\in(0,\tfrac{1}{2})$\,,\\ \,f(s)&for $H=\frac{1}{2}$\,,\\ \,C_{H}\int_{\mathbb{R}}\left|u-s\right|^{H-\frac{3}{2}}f(u)\mathop{}\!du&for $H\in(\tfrac{1}{2},1)$\,,\end{dcases*}\quad C_{H}=\frac{\sqrt{\Gamma(2H+1)\sin(\pi H)}}{2\Gamma\big(H-\frac{1}{2}\big)\cos\big[\frac{\pi}{2}\big(H-\frac{1}{2}\big)\big]}\,$, (S1)
where $\Gamma(\cdot)$ represents the gamma function. Choose $f(s)$ as the indicator function $\chi_{[a,b]}(s)$, which equals $1$ if $a\leq s\leq b$, 1-1 if $b\leq s\leq a$ (except when $s=a=b$), and 0 otherwise. We then integrate Eq. (S1) over the entire real line with respect to $\mathop{}\!dW(s)$. Setting $a=0$ and $b=t$, we obtain
$\int_{\mathbb{R}}M_{H}\chi_{[0,t]}(s)\mathop{}\!dW(s)=\frac{C_{H}}{\left|H-\frac{1}{2}\right|}\int_{\mathbb{R}}\Big[(t-s)^{H-\frac{1}{2}}\Theta(t-s)-(-s)^{H-\frac{1}{2}}\Theta(-s)\Big]\mathop{}\!dW(s)\,$, (S2)
where $\Theta(\cdot)$ denotes the Heaviside step function, defined as $\Theta(x)=1$ for $x>0$ and $\Theta(x)=0$ for $x\leq 0$. Equation (S2) matches, up to a normalisation factor, the original expression for $B_{H}(t)$ introduced by Mandelbrot and Van Ness [56, Eq. (2.1)].
@@
$X(t)$ $=X_{0}\exp\bigg(\int_{0}^{t}\alpha(s)\mathop{}\!ds+\int_{0}^{t}\beta(s)\mathop{}\!dB_{H}(s)-\frac{1}{2}\int_{\mathbb{R}}\big[M_{H}(\beta\circ\chi_{[0,t]})(s)\big]^{2}\mathop{}\!ds\bigg)\,$, (S4)
where the operator $M_{H}(\cdot)$ and the indicator function $\chi_{[a,b]}(\cdot)$ are as defined earlier. In the main text, we considered the special case where $\beta(s)=\beta$ is constant, and under that condition, the solution (S4) reduces to Eqs. (9) and (10).
Supplementary Text for Sec. IV
@@
In the main text, when simulating $C(T)$ as expressed in Eq. (7) based on our theoretical model, we introduced the approximation $\hat{C}(T)$ as given in Eq. (17). This approximation always yields a natural number for $T>0$, even in cases where $X_{i}\mathop{}\!\Delta t$ remains negligibly small throughout $i=1,\,\dots,\,n$, such that its summation over the entire period does not even reach 1. Nonetheless, Eq. (17) ensures that $\hat{C}=1$ even in such situations. In this respect, the effect is analogous to the $\text{citation}+1$’ prescription often used in the literature [35, 32, 24, 23], which facilitates the computation of the logarithm of citation counts even in cases with zero citations. We assume that the bias introduced by employing the ceiling function here is negligible and does not affect our main conclusions.
@@
Figure S2 shows simulated paths of $X(t)$ under $H=0.85$ and $\alpha(t)=1$. In contrast to the antipersistent case (Fig. 3), the trajectories here are notably smoother, with fewer pronounced fluctuations. In Fig. S2(c), $\{X_{i}\}_{i=1}^{n}$ follows a nearly steady trend and eventually declines to zero at time $T$. However, this outcome is not inevitable. Due to the persistent nature of the process, $\{X_{i}\}_{i=1}^{n}$ often continues along a certain trajectory, yet it could also move in the opposite direction and increase indefinitely. Such paths contribute to large values of $\hat{C}(T)$.
Figure S2 shows the results of Monte Carlo simulations with $N_{\mathrm{s}}=50{,}000$, conducted under the same conditions. Additionally, Fig. S3(a) presents the kernel density estimates of the distribution for $\beta=0.2$, $0.5$, $0.8$, $1.0$ and $1.5$. Compared with the antipersistent case (Fig. 3), the distribution of $C(T)$ now has a substantially fatter tail, a tendency that grows more pronounced with increasing $\beta$.
Moreover, the Q–Q plot in Fig. S3(b) deviates noticeably from a straight line, indicating that the distribution of $\hat{C}(T)$ no longer closely follows a log-normal form. This does not necessarily mean that a power law is the best fit. However, as shown in Fig. S3(c) and Table S1, where the upper 1% of $\hat{C}(T)$ is fitted with a shifted power law, the persistent scenario exhibits a stronger tendency to be well approximated by a shifted power law than the antipersistent case. Still, at extremely high citation counts, the data also deviate from the shifted power law.
Supplementary Text for Sec. V
@@
The data used in Fig. 1 and the empirical data analysis in Sec. V of the main text are the same as those employed in Ref. [23], with the raw data compiled in Ref. [55]. For further details on the data, please refer to these references.
@@
In the main text, the prospective approach is used (Sec. II.2). However, using the retrospective approach yields qualitatively similar results, again confirming the $t^{H}$ law trend. Indeed, when the $t^{H}$ law was first observed, the retrospective approach was applied [23, Appendix C, Suppl. Fig. S9].
@@
In the main text, the estimation of the Hurst parameter $H$ and volatility $\beta$ was carried out on the entire aggregate set of e-prints across six arXiv disciplines—Astrophysics (‘astro-ph’), Computer Science (‘comp-sci’), Condensed Matter Physics (‘cond-mat’), High Energy Physics (‘hep’), Mathematics (‘math’) and other physics (‘oth-phys’)—the same classification as in Refs. [23, 55]. Figure S6 shows the results of the same analysis, now broken down by discipline. From Fig. S6(a), it is clear that the memory effect is antipersistent in all fields, but a consistent historical trend indicates that High Energy Physics has the strongest antipersistent tendency, whereas Condensed Matter Physics has the weakest. Additionally, Fig. S6(b) reveals that volatility is especially high in High Energy Physics and Computer Science.
These observations likely reflect each discipline’s distinct characteristics, such as scientific discovery patterns, citation behaviours in research communities, and publishing norms—factors also discussed in the main text. Combined with earlier analyses of discipline-specific spatiotemporal aspects of citation evolution and with measures such as the ‘internal obsolescence rate’ and ‘retention rate’ reported in Ref. [23], these results provide useful insights not only from scientometric or bibliometric standpoints, but also within the broader ‘science of science’ framework [80].
@@
Typically, the Hurst parameter for fractional Brownian motion is estimated via methods such as the rescaled range analysis (R/S) method [72, 73] or subsequent methodological refinements. This approach involves calculating the range of partial sums of deviations for time series segments from their means, then normalising by the corresponding SD. However, it is not suitable in te current study. The reason is that the presence of an unknown, time-dependent deterministic nonlinear term $\alpha(t)$ disrupts self-similar scaling properties, thereby biasing R/S-based estimates.
Moreover, the arXiv dataset used in this study [55] records citation counts on a yearly basis, yielding at most a few dozen data points per e-print. Methods like R/S analysis [72, 73], detrended fluctuation analysis (DFA) [74, *Hu01, *Chen02], or alternative estimation techniques commonly employed in financial engineering [81, 57] usually require much longer time series. Despite recent advances in bibliometric data, collecting extended citation histories at finer temporal resolutions (e.g. monthly or daily) remains practically challenging. Where monthly-level citation records exist, they often involve data imputation: for instance, if the exact publication month is missing, the record may default to January. Similarly, daily-level data can default to the first of the month if only the month is known. Hence, careful preprocessing is needed to handle missing values or exclude certain entries, making the construction of precise time series difficult at monthly or daily resolutions.
Given these constraints, this study employed a heuristic estimation of the Hurst parameter based on the newly discovered $t^{H}$ law. Although the resulting estimates are approximate, their qualitative features—such as the evidence of antipersistent behaviour—are statistically robust within the limits of the applied tests. It should nevertheless be emphasised that these results are derived from a simplified estimation procedure and from data with a limited number of points, and must therefore be interpreted with due regard to the associated biases and limitations.
@@
Now consider a hypothetical world in which all publishing constraints—whether resource-based, technological, or conventional—are entirely removed. In this setting, researchers would have optimal access to literature search, knowledge acquisition, and reference management. Citing older publications would be actively encouraged, with no practical limit on the number of references included. Under such conditions, the typical citation history curve, when aggregated across numerous publications, might no longer follow a jump-and-decay pattern but instead develop a heavy-tailed form over time or even evolve into a monotonically increasing function. In this scenario, the distribution of the time-integrated attention measure would likely be better represented by a power law or another even heavier-tailed distribution, rather than a log-normal distribution. Alternatively, it is possible that such an attention-dynamics paradigm already exists in certain domains, even if it has not yet been widely recognised.
@@
Figure S4: Relationship between the standard deviation (SD) of the logarithm of citation counts acquired each year up to February 2020 and the number of years ($t$) elapsed since posting, shown for e-prints published on arXiv between 1995 and 2006. The data used are the same as in Ref. [55].
Figure S5: Estimated values of (a) $\hat{H}$ and (b) $\hat{\beta}$, along with their standard errors for e-prints published on arXiv from 1995 to 2006, presented for each publication year (numerical data are provided in Table S2). The data used are the same as in Ref. [55].
Figure S6: Estimated values of (a) $\hat{H}$ and (b) $\hat{\beta}$ for e-prints published on arXiv from 1999 to 2010, shown for six different arXiv disciplines—Astrophysics (‘astro-ph’), Computer Science (‘comp-sci’), Condensed Matter Physics (‘cond-mat’), High Energy Physics (‘hep’), Mathematics (‘math’) and other physics (‘oth-phys’). The data used are the same as in Ref. [55].